### v3

# UniHR: Hierarchical Representation Learning for Unified Knowledge Graph Link Prediction

Abstract
Beyond-triple fact representations including hyper-relational facts with auxiliary key-value pairs, temporal facts with additional timestamps, and nested facts implying relationships between facts, are gaining significant attention. However, constrained by complex fact representation forms, existing link prediction models for beyond-triple facts have difficulty achieving hierarchical fact modeling and generalizing the modules for one specific facts to other fact types. To overcome this limitation, we propose a Unified Hierarchical Representation learning framework (UniHR) for unified knowledge graph link prediction. It consists of a unified Hierarchical Data Representation (HiDR) module and a unified Hierarchical Structure Learning (HiSL) module as graph encoder. The HiDR module unifies hyper-relational KGs, temporal KGs, and nested factual KGs into triple-based representations. Then HiSL incorporates intra-fact and inter-fact message passing, focusing on enhancing the semantic information within individual facts and enriching the structural information between facts. Empirical results demonstrate the effectiveness of UniHR and highlight the strong potential of unified representations. Code and data are available at https://github.com/Lza12a/UniHR.
Keywords:
Knowledge graph representation learning Beyond-triple knowledge graph Link prediction.
1 Introduction
Large-scale knowledge graphs (KGs) such as Freebase [3] and Wikidata [33] have been widely applied in many areas like question answering [17] and natural language processing [2]. However, the presence of missing facts within these KGs inevitably limit their applications. Therefore, the link prediction task has been introduced to predict missing elements within factual data. Current link prediction methods mainly focus on the fact in the form of triple e.g., $(head\,entity,relation,$
$tail\,entity)$.
Figure 1: Comparison between the existing link prediction methods for specific beyond-triple knowledge graphs and our unified methods.
Despite the simplicity and unity of triple-based representation, it is difficult to adequately express complex facts, such as “Oppenheimer is educated at Harvard University for a bachelor degree in chemistry" shown in Figure 1. Therefore, existing researches [34, 37, 38] contribute to focusing on semantically richer facts. Figure 1 illustrates three specific types of facts: hyper-relational fact ((Oppenheimer, educated at, Harvard University), degree: bachelor, major: chemistry), temporal fact (Oppenheimer, honored with, Fermi Prize, 1963), nested fact ((Oppenheimer, born in, New York), imply, (Oppenheimer, nationality, The United States)). These forms of facts allow for expression of complex semantics and revelation of relationships between facts, extending beyond the triple-based representation. Thus in recent years, Hyper-relational KGs (HKG) [6], Temporal KGs (TKG) [40], and Nested factual KGs (NKG) [37] attract wide research interests.
Recent studies have demonstrated the effectiveness of various embedding strategies for these beyond-triple representations [36]. However, these methods are hard to achieve hierarchical fact modeling and usually designed for specific representation forms, e.g., StarE [12] customizes graph neural network to enhance inter-fact structure information between hyper-relational facts. For NKGs, BiVE [7] bridges intra-fact semantics between two levels of facts through a simple MLP. Moreover, GeomE+ [40] et al. TKG methods employ time-aware scoring functions to capture intra-fact collaborative information between factual semantics and temporal information in the timestamp. Although these methods perform well on specific types of facts, it is evident that such customized methods are difficult to generalize to other types of KGs for intra-fact structure and inter-fact semantic enhancements. Therefore, establishing a unified hierarchical representation learning method for multiple types of KGs is worth to investigate.
To overcome the challenges mentioned above, we propose a Unified Hierarchical Representation learning method (UniHR), which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module as the graph encoder. HiDR module standardizes hyper-relational facts, nested factual facts, and temporal facts into the form of triples without loss of information. Furthermore, HiSL module captures local semantic information during intra-fact message passing and then utilizes inter-fact message passing to enrich the global structure information to obtain better node embeddings based on HiDR form. Finally, the updated embeddings are fed into decoders for link prediction. Empirical results demonstrate the UniHR achieves state-of-the-art or competitive performance on most metrics and highlight the strong potential of unified representations. Our contributions can be summarized as follows.
  • 1.
    We emphasize the value of investigating unified KG representation method, including unified symbolic representation and unfied representation learning method for different KGs.
  • 2.
    To our knowledge, we propose the first unified KG representation learning framework UniHR, across different types of KGs, including a hierarchical data representation module and a hierarchical structure learning module.
  • 3.
    We conduct link prediction experiments on 9 datasets across 5 types of KGs. Compared to methods designed for one kind of KG, UniHR achieves the best or competitive results, verifying strong generalization capability.
2 Preliminaries
In this section, we introduce the definition of four types of existing knowledge graphs (KGs): triple-based KG, hyper-relational KG, nested factual KG and temporal KG, along with link prediction tasks on these types of KGs.
Link Prediction on Triple-based KG.
A triple-based KG $\mathcal{G}_{KG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ represents facts as triples, denoted as =$\mathcal{F}\,\text{=}\{\left(h,r,t\right)|h,t\in\mathcal{V},$ $r\in\mathcal{R}\}$, where $\mathcal{V}$ is the set of entities and $\mathcal{R}$ is the set of relations. The link prediction on triple-based KGs involves answering a query $\left(h,r,?\right)$ or (?$,r,t\right)$, where the missing element ‘????’ is an entity in $\mathcal{V}$.
Link Prediction on Hyper-relational KG.
A hyper-relational KG (HKG) $\mathcal{G}_{HKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ consists of hyper-relational facts, abbreviated as H-Facts, denoted as =$\mathcal{F}\,\text{=}\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})$ $|\,h,t,v_{i}\in\mathcal{V},r,k_{i}\in\mathcal{R}\}$. Typically, we refer to $\left(h,r,t\right)$ as the main triple in the H-Fact and $\left\{\left(k_{i}\text{:}\,v_{i}\right)\right\}_{i=1}^{m}$ as $m$ auxiliary key-value pairs. Similar to link prediction on triple-based KGs, the link prediction on HKGs aims to predict entities in the main triple or the key-value pairs. Symbolically, the aim is to predict the missing element, denoted as ‘?’ for queries $((h,r,t),(k_{1}\text{:}\,v_{1}),\ldots(k_{i}\text{:}\,?)),$ ((?$,r,t),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$ or $((h,r,?),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$.
Link Prediction on Nested Factual KG.
A nested factual KG (NKG) can be represented as $\mathcal{G}_{NKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\hat{\mathcal{R}},\hat{\mathcal{F}}\}$, which is composed of two levels of facts, called atomic facts and nested facts. =$\mathcal{F}\,\text{=}\left\{\left(h,r,t\right)|h,t\in\mathcal{V},r\in\mathcal{R}\right\}$ is the set of atomic facts, where $\mathcal{V}$ is a set of atomic entities and $\mathcal{R}$ is a set of atomic relations. ^=$\hat{\mathcal{F}}\,\text{=}\,\{\left(\mathcal{F}_{i},\hat{r},\mathcal{F}_{j}\right)|$ $\mathcal{F}_{i},\mathcal{F}_{j}\in\mathcal{F},\hat{r}\in\hat{\mathcal{R}}\}$ is the set of nested facts, where $\hat{\mathcal{R}}$ is the set of nested relations. The link prediction on the NKGs is performed on the atomic facts or nested facts. We refer to the link prediction on atomic facts as Base Link Prediction, and the link prediction on nested facts as Triple Prediction. For base link prediction, given a query $\left(h,r,?\right)$ or (?$,r,t\right)$, the aim is to predict the missing atomic entity ‘????’ from $\mathcal{V}$. For triple prediction, given a query (?$,\hat{r},\mathcal{F}_{j}\right)$ or $\left(\mathcal{F}_{i},\hat{r},?\right)$, the aim is to predict the missing atomic fact ‘????’ from $\mathcal{F}$.
Link Prediction on Temporal KG.
A temporal KG (TKG) $\mathcal{G}_{TKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\mathcal{T}\}$ is composed of quadruple-based facts, which can be represented as =$\mathcal{F}\,\text{=}\,\{(h,r,t,$
$[\tau_{b},\tau_{e}])|h,t\in\mathcal{V},r\in\mathcal{R},\tau_{b},\tau_{e}\in\mathcal{T}\}$, where $\tau_{b}$ is the begin time, $\tau_{e}$ is the end time, $\mathcal{V}$ is the set of entities, $\mathcal{R}$ is the set of relations and $\mathcal{T}$ is the set of timestamps. The link prediction on TKGs aims to predict missing entities ‘????’ in $\mathcal{V}$ for two types of queries (?$,r,t,[\tau_{b},\tau_{e}]\right)$ or $\left(h,r,?,[\tau_{b},\tau_{e}]\right)$.
3 Related Works
Link Prediction on Hyper-relational Knowledge Graph.
Earlier HKG representation learning methods e.g., m-TransH [35], RAE [44] have generalized the triple-based approach to HKG and loosely represent the combinations of key-value pairs. Galkin et al. first customize StarE [12] based on CompGCN [31] for H-Facts to capture global structure information between H-Facts in the message passing stage, and achieves impressive results, demonstrating that the structure information of the graph in HKGs is also important. GRAN [15] introduces edge-aware bias into the vanilla transformer attention [32], while HyNT [6] designs a qualifier encoder for HKG. They both focus on intra-fact semantic dependencies but ignore the global structure information. In order to fully capture inter-fact and intra-fact information, HAHE [23] first adopts dual graph attention and edge-aware bias in the transformer attention layer for hierarchical modeling and achieves significant performance improvement. We consider the hierarchical attention mechanism of HAHE a good beginning, but it is customized for HKG and is difficult to generalize to other facts.
Link Prediction on Nested Factual Knowledge Graph.
Chung et al. [7] first introduced nested facts to imply relationships between facts. They also propose BiVE which bridges semantics between atomic facts and fact nodes in the encoding phase via a simple MLP and scores both atomic facts and nested facts using the quaternion-based KGE scoring functions like QuatE [45] or BiQUE [16]. Based on BiVE, NestE [37] represents the fact nodes as a $1\times 3$ embedding matrix and the nested relations as a $3\times 3$ matrix to avoid information loss, embedding them into hyperplanes with different dimensions. These methods only capture the intra-fact semantic information between atomic facts and nested facts. Meanwhile, due to the complexity of this representation, common triple-based GNNs have difficulty in message passing between atomic fact and nested fact for capturing inter-fact global information.
Link Prediction on Temporal Knowledge Graph.
Recent studies in temporal knowledge graph representation learning have focused on designing special time-aware scoring functions for inter-fact semantic modeling. Models such as TTransE [21], HyTE [8], TeRo [39], and TGeomE+ [40] incorporate temporal-aware module into the KGE score function in various ways. However, existing models seldom regard timestamps as node and directly utilize inter-fact message passing mechanism to perceive time information for enhancing entity and relation embeddings. Therefore, we believe it is a direction worth exploring.
4 Methodology
In this section, we introduce our method, a Unified Hierarchical Representation learning framework (UniHR), which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module. Our workflow can be divided into the following three steps: 1) Given a KG $\mathcal{G}$ of any type, we represent it into $\mathcal{G}^{HiDR}$ under the HiDR form. 2) The $\mathcal{G}^{HiDR}$ will be encoded by HiSL module to enhance the semantic information within individual facts and structural information between facts on the whole graph. 3) In the phase of decoding, the updated embeddings of nodes and edges are fed into transformer decoders to obtain the plausibility score of facts.
Figure 2: Diverse facts are translated into HiDR form.
4.1 Hierarchical Data Representation
To overcome the differences in the representation of multiple types KGs, we introduce a Hierarchical Data Representation module, abbreviated as HiDR. Different from labelled RDF representation [1], we constrain “triple" to be considered as the basic units of HiDR form, then HiDR could continuous benefit from the model developments of triple-based KGs, which is the most active area about link prediction over KGs.
As shown in Fig. 2, in order to ensure comprehensive representation of facts, we introduce three hierarchical types of nodes and three connected relations in HiDR. Firstly, we denote original entities within three types of KGs as atomic nodes and complement fact nodes for HKGs and TKGs lacking a designated fact node. To facilitate the interaction between fact nodes and relations explicitly, we incorporate relation nodes into the graph, represented as $e_{r}$ for each $r$. These relation nodes are derived from transforming the relation edges in the original KG. It is important to facilitate direct access of fact nodes to the relevant atomic nodes during message passing process. To achieve this, we introduce three connected relations: has relation, has head entity and has tail entity, which establish directly connections between atomic nodes and fact nodes. Ultimately, we denote the (main) triple $\left(h,r,t\right)$ in original fact as three connected facts: $\left(f,has\,relation,e_{r}\right),\left(f,has\,head\,entity,h\right),$ $(f,has\,tail\,entity,t)$, and an atomic fact $(h,r,t)$, where $f$ is fact node. Formally, the definition of HiDR form is as follows:
Definition 1
Hierarchical Data Representation (HiDR): A KG represented as the HiDR form is denoted as $\mathcal{G}^{HiDR}\,\text{=}\,\{\mathcal{V}^{HiDR},\mathcal{R}^{HiDR},\mathcal{F}^{HiDR}\},$ where $\mathcal{V}^{HiDR}\,\text{=}\,$
$\mathcal{V}_{a}\cup\mathcal{V}_{r}\cup\mathcal{V}_{f}$ is a joint set of atomic node set ($\mathcal{V}_{a}$), relation node set ($\mathcal{V}_{r}$), fact node set ($\mathcal{V}_{f}$). $\mathcal{R}^{HiDR}\,\text{=}\,\mathcal{R}_{a}\cup\mathcal{R}_{n}\cup\mathcal{R}_{c}$ is a joint set of atomic relation set ($\mathcal{R}_{a}$), nested relation set ($\mathcal{R}_{n}$), connected relation set $\mathcal{R}_{c}\,\text{=}\{has\,relation,has\,head\,entity,has\,$
$tail\,entity\}$. The fact set $\mathcal{F}^{HiDR}\,\text{=}\,\mathcal{F}_{a}\cup\mathcal{F}_{c}\cup\mathcal{F}_{n}$ is jointly composed of three types of triple-based facts: atomic facts ($\mathcal{F}_{a}$), connected facts ($\mathcal{F}_{c}$) and nested facts ($\mathcal{F}_{n}$), where $\mathcal{F}_{a}\text{=}\,\{(v_{1},r,$ $v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{a},r\in\mathcal{R}_{a}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1}\in\mathcal{V}_{f},r\in\mathcal{R}_{c},v_{2}\in\mathcal{V}_{a}\}$, $\mathcal{F}_{n}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{f},r\in\mathcal{R}_{n}\}$.
Next, we introduce how to transform different types of KGs into HiDR form.
For hyper-relational knowledge graphs, we regard key-value pairs as complementary information for facts. Thus, we translate the H-Facts $\mathcal{F}_{HKG}\,\text{=}$
$\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m}\}$ into the HiDR form that $\mathcal{G}^{HiDR}_{HKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{HKG}\}$ following the definition, where $\mathcal{F}_{c}\,\text{=}\,\{(f,has\,relation,e_{r}),(f,has$ $\,head\,entity,h),(f,has$
$tail\,entity,t),(f,k_{1},v_{1}),\ldots,(f,k_{m},v_{m})\}$, $\mathcal{F}_{a}\,\text{=}\,\{(h,r,t)\,|\,((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})\in\mathcal{F}_{HKG})\}$ and $\mathcal{F}_{n}\text{=}\,\varnothing$.
For nested factual knowledge graphs, HiDR can naturally represent hierarchical facts, so we translate the atomic facts $\mathcal{F}_{NKG}\,\text{=}\,\{(h_{i},r_{i},t_{i})\}$ and the nested facts ^$\hat{\mathcal{F}}_{NKG}\,\text{=}\,\{((h_{1},r_{1},t_{1}),R,(h_{2},$ $r_{2},t_{2}))|(h_{i},r_{i},t_{i})\in\mathcal{F}_{NKG}\}$ into the form of HiDR that $\mathcal{G}^{HiDR}_{NKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{NKG}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\,\{(h_{i},r_{i},t_{i})\,|\,(h_{i},r_{i},t_{i})\in$ $\mathcal{F}_{NKG}\}$, $\mathcal{F}_{c}\,\text{=}\{(f_{i},has\,head\,entity,h_{i}),(f_{i},has\,tail$
$entity,t_{i}),(f_{i},has\,relation,e_{r_{i}})|f_{i}\,\text{=}\,(h_{i},r_{i},$ $t_{i})\in\mathcal{F}_{NKG}\}$ and $\mathcal{F}_{n}\text{=}\{(f_{1},R,f_{2})|f_{i}\in\mathcal{F}_{NKG}\}$.
For temporal knowledge graphs, we regard the TKG as a special HKG, and convert timestamps to auxiliary key-value pairs in HKGs by adding two special atomic relations: begin and end, regarding timestamps as special numerical atomic nodes. Thus, we firstly translate the temporal facts in TKGs $\mathcal{F}_{TKG}\text{=}\,\{(h,r,t,[\tau_{b},\tau_{e}])\}$ into H-Facts form $\mathcal{F}_{TKG}^{HKG}\text{=}\,\{(h,r,t,begin\text{:}\tau_{b},end\text{:}\tau_{e})\}$. Then according to the previous transformation in HKG, it can be translated into the HiDR form that $\mathcal{G}_{TKG}^{HiDR}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}_{TKG}^{HiDR}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\{(h,r,t)\,|\,(h,r,t,begin\text{:}\,\tau_{b},end\text{:}$ $\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(f,has\,relation,e_{r}),(f,$ $has\,head\,entity,h),\,(f,has\,tail\,entity,t),\,(f,$ $begin,\tau_{b}),(f,end,\tau_{e})\,|\,f\,\text{=}\,(h,r,t,$
$begin\text{:}\,\tau_{b},end\text{:}\,\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$ and $\mathcal{F}_{n}\,\text{=}\,\varnothing$.
In summary, we could convert all above KGs into the HiDR form, and preserve the semantics in the original KGs without loss of information.
Figure 3: HiSL module for intra-fact and inter-fact message passing.
4.2 Hierarchical Structure Learning
It’s evident that HiDR form introduces many additional relation nodes and fact nodes. To avoid significantly increasing the model’s training parameters while fully capturing the hierarchy of HiDR form, we design a Hierarchical Structure Learning module, abbreviated as HiSL shown in Fig. 3.
Representation Initialization.
We first initialize the embedding matrices $\mathbf{H}_{a}\in\mathbb{R}^{|\mathcal{V}_{a}|\times d}$ and $\mathbf{E}\in\mathbb{R}^{|\mathcal{R}^{HiDR}|\times d}$ for atomic nodes and all relation edges. Then we also initialize the embedding of relation node $\mathbf{H}_{r}\in\mathbb{R}^{|\mathcal{V}_{r}|\times d}$, which can be transformed from the relation edge $r$ with a projection matrix $\mathbf{W_{r}}\in\mathbb{R}^{d\times d}$: $\mathbf{H}_{r}=\mathbf{E_{a}\cdot W_{r}}$, where $\mathbf{E_{a}}\subseteq\mathbf{E}$ is the atomic relation embeddings. Then we initialize the fact node embeddings $\mathbf{H}_{f}$ to explicitly capture key information within facts by utilizing the embedding of (main) triple:
$\mathbf{h}_{f}=f_{m}([\mathbf{h}_{h};\mathbf{h}_{r};\mathbf{h}_{t}]),$ (1)
where $(h,r,t)\in\mathcal{F}_{a}$, [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] is the concatenation operation, $\mathbf{h}_{h},\mathbf{h}_{t}\subseteq\mathbf{H}_{a},\mathbf{h}_{r}\subseteq\mathbf{H}_{r}$ denote (main) triple embedding and $f_{m}\text{:}\;\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP. Therefore, the initialization of relation nodes and fact nodes is sufficiently parameter-efficient.
For numerical atomic nodes, namely timestamps in temporal knowledge graphs, we encode the timestamp $\tau$ into an embedding with Time2Vec [18]:
$\mathbf{h}_{\tau}=\omega_{p}\sin\left(f_{p}(\tau)\right)+f_{np}(\tau),$ (2)
where $f_{p}\text{:}\;\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP as periodic function, $f_{np}\text{:}\;\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP as non-periodic function, and $\omega_{p}\in\mathbb{R}^{1}$ is a learnable parameter for scaling the periodic features.
Intra-fact Message Passing.
In this stage, massage passing is conducted for fact nodes. Given a fact node $f_{k}\in\mathcal{V}_{f}$, we construct its constituent elements, i.e., one-hop neighbors, as the node set $\mathcal{V}_{k}\,\text{=}\,\{v\in\mathcal{N}_{f_{k}}\,|\,v\in\mathcal{V}_{a}\cup\mathcal{V}_{r}\}$, where $\mathcal{N}_{f_{k}}$ is the set of one-hop neighbors of fact node $f_{k}$. Then we retain the edges directly connected to fact node $f_{k}$, thereby constructing a subgraph $\mathcal{G}_{k}\,\text{=}\,\{\mathcal{V}_{k},\mathcal{R}_{k},\mathcal{F}_{k}\}\subseteq\mathcal{G}^{HiDR}$. For this subgraph, we employ the graph attention network [5] to aggregate local information, computing the attention score $\alpha_{i,j}$ between node $i\in\mathcal{V}_{k}$ and its neighbor $j$. The formula for calculating $\alpha_{i,j}$ in the $l$-th layer is as follows:
$\alpha_{i,j}^{l}=\frac{\exp\left(\mathbf{W}^{l}\left(\sigma\left(\mathbf{W}_{in}^{l}\mathbf{h}_{i}^{l}+\mathbf{W}_{out}^{l}\mathbf{h}_{j}^{l}\right)\right)\right)}{\sum\limits_{j^{\prime}\in\mathcal{N}_{i}}\exp\left(\mathbf{W}^{l}\left(\sigma\left(\mathbf{W}_{in}^{l}\mathbf{h}_{i}^{l}+\mathbf{W}_{out}^{l}\mathbf{h}_{j^{\prime}}^{l}\right)\right)\right)},$
where $\mathbf{h}_{i}^{l},\mathbf{h}_{j}^{l}\in\mathbb{R}^{d}$ represent the embeddings of node $i$ and its neighbor $j$ in $l$-th layer. And there are three learnable weight matrices $\mathbf{W}_{in}^{l},\mathbf{W}_{out}^{l}\in\mathbb{R}^{d\times d}$ and $\mathbf{W}^{l}\in\mathbb{R}^{d}$. We choose LeakyReLU as activation function $\sigma$. Then, the updated node embeddings are obtained by aggregating the information of neighbors according to the attention scores:
$\mathbf{h}_{i}^{l}=\mathbf{h}_{i}^{l}+\sum\limits_{j\in\mathcal{N}_{i}}\alpha_{i,j}^{l}\cdot\mathbf{W}_{out}^{l}\mathbf{h}_{j}^{l}.$ (4)
Inter-fact Message Passing.
At this stage, message passing is conducted on the whole graph $\mathcal{G}^{HiDR}$. Similar to previous work [31], we use a non-parametric aggregation operator $\phi\left(\cdot\right)\text{:}\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to obtain messages from neighbouring nodes and edges. We employ the circular-correlation operator, defined as:
$\phi\left(\mathbf{h}_{j},\mathbf{e}_{r}\right)=\mathbf{h}_{j}\star\mathbf{e}_{r}=\mathbf{F}^{-1}\left(\left(\mathbf{F}\mathbf{h}_{j}\right)\odot\overline{\left(\mathbf{F}\mathbf{e}_{r}\right)}\right),$ (5)
where $\mathbf{F}$ and $\mathbf{F}^{-1}$ denote the discrete fourier transform (DFT) matrix and its inverse matrix, the $\odot$ is the element-wise (Hadamard) product. Furthermore, in order to fully capture the heterogeneity of the graph, we classify edges along two dimensions: $\lambda(r)\in\left\{forward,reverse\right\}$ and $\tau(r)\in\{connected\,\,relation,atomic\,\,$
$relation,nested\,relation\}$ and adopt two relation-type specific learnable parameters $\mathbf{W}_{\lambda\left(r\right)}\in\mathbb{R}^{d\times d}$ and $\omega_{\tau\left(r\right)}\in$ $\mathbb{R}^{1}$ for more fine-grained aggregation:
$\mathbf{h}_{i}^{l+1}=\sum\limits_{(r,j)\in\mathcal{N}(i)}{\sigma\left(\omega_{\tau\left(r\right)}^{l}\right)\mathbf{W}_{{\lambda}\left(r\right)}^{{l}}\phi\left(\mathbf{h}_{j}^{l},\mathbf{e}_{r}^{l}\right)}+\mathbf{W}_{self}^{l}\mathbf{h}_{i}^{l},$ (6)
$\mathbf{e}_{r}^{l+1}=\mathbf{W}_{rel}^{l}\mathbf{e}_{r}^{l},$ (7)
where $\mathbf{W}^{l}_{self},\mathbf{W}^{l}_{rel}\in\mathbb{R}^{d\times d}$, $\sigma$ is a sigmoid activation function and $\mathcal{N}\left(i\right)$ is a set of immediate neighbors of $i$ for its outgoing edges $r$. We utilize $\phi\left(\cdot\right)$ to combine the information from edge $r$ and node $j$, and then passes it to node $i$ for update.
Through Intra-fact and Inter-fact two-stage message passing, nodes can fully capture both local semantic and global structural information. Moreover, its number of training parameters does not increase with the scale of the graph, thereby effectively adapting to the HiDR form.
4.3 Link Prediction Decoder
Since the query varies across different settings, we use the transformer [32] as the decoder with the mask pattern. Specifically, we convert the updated node and edge embeddings into a sequence of fact embeddings, mask the elements to be predicted in facts with the $\bm{\left[M\right]}$ token as the input to the transformer. Finally, we obtain the embedding of output $\bm{\left[M\right]}$ in the last layer to measure the plausibility of the fact, denoted as $\mathbf{h}_{pre}$, and calculate the probability distribution of candidates, followed by training it using the cross-entropy loss:
$P=\,{\rm Softmax}\left(f\left(\mathbf{h}_{pre}\right)[\mathbf{E};\mathbf{H}]^{\top}\right),$ (8)
$\mathcal{L}=\sum_{t=0}^{\left|\mathcal{R}\right|+\left|\mathcal{V}\right|}{y_{t}\log P_{t}},$ (9)
where $P\in\mathbb{R}^{\left|\mathcal{R}\right|+\left|\mathcal{V}\right|}$ represents the confidence scores of all candidates, $f\text{:}\;\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP, and $[\mathbf{E};\mathbf{H}]\in\mathbb{R}^{(\left|\mathcal{R}\right|+\left|\mathcal{V}\right|)\times d}$ is the embedding matrix of all candidate edges or nodes. The $P_{t}$ and $y_{t}$ are probability and ground truth of the $t$-th candidate. The final loss function $\mathcal{L}$ includes both node loss and edge loss during the predictions.
5 Experiment
5.1 Experiment Settings
Datasets.
For link prediction on HKGs, we select three benchmark datasets: WikiPeople [15] and WD50K [12]. For link prediction on NKGs, we choose FBH, FBHE and DBHE constructed by [7]. For link prediction on TKGs, we use wikidata12k [8], a subset of wikidata [33]. To further evaluate the potential of the unified representation, we further introduce the hyper-relational TKG datasets WIKI-hy and YAGO-hy [10]. The detailed statics are given in Table 1.
Table 1: The statistics of diverse knowledge graphs dataset, where “with Q(%)" and “Arity" column respectively denote the number of facts with auxiliary key-value pairs and the range of arity of hyper-relational facts, “N-Fact” is the number of nested fact, “N-Rela” is the number of nested relation, the “AF(%)" column denotes the number of atomic facts in nested facts.
Dataset Fact Entities Rela Train Valid Test with Q(%) Arity N-Fact N-Rela AF(%) Period
Hyper-relational Knowledge Graph
WikiPeople 369866 34825 178 294439 37715 37712 9482(2.6%) 2-7 - - - -
WD50K 236507 47155 531 166435 23913 46159 32167(13.6%) 2-67 - - - -
Nested Factual Knowledge Graph
FBH 310116 14541 237 248094 31011 31011 - - 27062 6 33157 -
FBHE 310116 14541 237 248094 31011 31011 - - 34941 10 33719 -
DBHE 68296 12440 87 54636 6830 6830 - - 6717 8 8206 -
Temporal Knowledge Graph
wikidata12k 40621 12554 24 32497 4062 4062 - - - - - 19-2020
Hyper-relational Temporal Knowledge Graph
WiKi-hy 139078 16634 147 111252 13900 13926 13335(9.59%) 2-8 - - - 1513-2020
YAGO-hy 73143 16167 54 51193 10973 10977 5107(6.98%) 2-5 - - - 0-187
Multiple types of Knowledge Graph
wikimix 409566 43832 184 326936 41777 3525/37328 9098(2.2%) 2-7 - - - 19-2020
Evaluation Metric.
We conduct link prediction across multiple settings, evaluating performance based on the rank of predicted facts. We use the MR (Mean Rank), MRR (Mean Reciprocal Rank) and Hits@K (K=1,3,10) as our evaluation metrics. We abbreviate ‘Hits@K’ as ‘H@K’ and employ filtering settings [4] during the evaluation to eliminate existing facts in the dataset. It is worth noting that for the query $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ in HKGs, there are two evaluation filtering settings in existing models: one that filters out facts satisfying $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ and another that filters out facts satisfying only $(h,r,?)$ in the training set. Similarly, the difference in filtering settings of TKG occurs in timestamp. In this paper, we adopt the strict filtering setting of the former. To ensure fair comparison, for HKG we utilize the results of HyNT [6] with the same settings as ours. For TKG, we thoroughly review the original code of our baselines and reproduce the results of some methods.
Baselines.
For link prediction on HKG, we compare our UniHR with NaLP [15], tNaLP [15], RAM [22], HINGE [27], NeuInfer [14], StarE [12], HyTransformer [43], GRAN [34], HyNT [6] and HAHE [23]. For link prediction on NKG, QuatE [45], BiQUE [16], Neural-LP [42], DRUM [28], AnyBURL [24], BiVE [7] and NestE [37] are chosen as baselines. Among them, BiVE and NestE are especially designed for NKG. We compare against following TKG link prediction methods: ComplEx-N3 [20], HyTE [8], TA-DistMult [13], ATiSE [38], TeRo [39], TGeomE+ [40], HGE [26] and ECEformer [11].
Implementation details.
All experiments are conducted on a single Nvidia 80G A800 GPU and implemented with PyTorch. For base link prediction on NKGs, we also use augmented triples from [7] for training to ensure fairness. For triple prediction, due to the small size of training set, we conduct training based on fixed embeddings of entities obtained from the base link prediction and set $\omega_{nested\;relation}\text{=}\ 0$ to prevent overfitting. We employ AdamW [19] optimizer. Hyperparameter details can be found in Appendix C.
5.2 Main Results
Table 2: Results of link prediction on HKG datasets. Results are reported by us, and the others are taken from [6]. Best results are bold, while the second are underlined.
WikiPeople WD50K
Model subject/object all entities subject/object all entities
MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
NaLP [15] 0.356 0.271 0.499 0.360 0.275 0.503 0.230 0.170 0.347 0.251 0.187 0.375
tNaLP [15] 0.358 0.288 0.486 0.361 0.290 0.490 0.221 0.163 0.331 0.243 0.182 0.360
RAM [22] 0.459 0.384 0.584 0.461 0.386 0.585 0.276 0.210 0.399 0.296 0.232 0.416
HINGE [27] 0.393 0.309 0.547 0.395 0.311 0.549 0.264 0.187 0.410 0.277 0.200 0.424
NeuInfer [14] 0.357 0.247 0.533 0.357 0.248 0.532 0.220 0.154 0.347 0.225 0.158 0.355
StarE [12] 0.458 0.364 0.611 - - - 0.309 0.234 0.452 - - -
HyTransformer [43] 0.465 0.395 0.598 - - - 0.308 0.239 0.452 - - -
GRAN [34] 0.477 0.408 0.596 0.480 0.411 0.599 0.329 0.259 0.465 0.363 0.294 0.493
HyNT [6] 0.482 0.415 0.602 0.481 0.414 0.603 0.333 0.259 0.474 0.360 0.287 0.500
HAHE [23] 0.498 0.426 0.610 0.497 0.429 0.614 0.343 0.270 0.484 0.378 0.306 0.515
UniHR 0.491 0.417 0.618 0.493 0.420 0.621 0.348 0.278 0.482 0.382 0.313 0.513
Link Prediction on HKG.
We compare our method with previous methods on the WD50K and WikiPeople datasets shown in Table 2. We can observe that the models based on transformers (i.e., StarE, GRAN, HyNT, HAHE) demonstrate significantly better performance compared to other models. We attribute this to the transformer’s superior ability to capture long distance dependencies within H-Facts. Among these methods, it can be seen that our proposed UniHR achieves competitive results with the state-of-the-art method HAHE, which means our method effectively captures hierarchical fact information. Compared to the GNN-based method StarE, we achieve improvements of 3.9 points (12.6%) in MRR, 4.4 points (18.8%) in Hits@1 and 3.0 points (6.6%) in Hits@10 on WD50K. This indicates that the performance of StarE’s customized graph neural network is limited by its inability to flexibly capture key-value pair information. Moreover, since the embeddings for newly added fact nodes and relation nodes are computed from atomic facts through simple MLPs, so our training parameters do not significantly increase.
Table 3: Results of base link prediction (left) and triple prediction (right). All baselines’ results are taken from [37]. We pick the best variants of BiVE [7] and NestE [37].
FBH DBHE FBH FBHE DBHE
Model MRR H@10 MRR H@10 MR MRR H@10 MR MRR H@10 MR MRR H@10
Base link prediction Triple prediction
QuatE [45] 0.354 0.581 0.264 0.440 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
BiQUE [16] 0.356 0.583 0.274 0.446 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
Neural-LP [42] 0.315 0.486 0.233 0.357 115016.6 0.070 0.073 90000.4 0.238 0.274 21130.5 0.170 0.209
DRUM [12] 0.317 0.490 0.237 0.359 115016.6 0.069 0.073 90000.3 0.261 0.274 21130.5 0.166 0.209
AnyBURL [24] 0.310 0.526 0.220 0.364 108079.6 0.096 0.108 83136.8 0.191 0.252 20530.8 0.177 0.214
BiVE [7] 0.370 0.607 0.274 0.422 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
NestE [37] 0.371 0.608 0.289 0.443 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
UniHR 0.401 0.619 0.296 0.448 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
Link Prediction on NKG.
Our experiments on the NKGs consist of two tasks: base link prediction and triple prediction. From the results in Table 3, we can see that our proposed UniHR obtains competitive results as the first method to capture global structural information of NKGs. For base link prediction task on triple-based KGs, UniHR achieves considerable improvements. Of particular note, the MRR of FBHE increases by 8.1%.
For triple prediction, we perform best on FBH and DBHE datasets, especially obtaining an improvement of 2.4 points in MRR on FBH, and achieve the second-best performance on FBHE, which suggests that structural information is also valuable for NKG and UniHR can effectively capture the heterogeneity of NKG to enhance node embeddings. In particular, as a unified method, we do not use the customized decoder for triples, while previous state-of-the-art methods do. We will further illustrate the effectiveness of UniHR equipped with other decoders in Appendix B.
Table 4: Results of link prediction on wikidata12k. Results are reported by us, and others are taken from original papers.
Model wikidata12k
MRR H@1 H@3 H@10
ComplEx-N3 [20] 0.248 0.143 - 0.489
HyTE [8] 0.253 0.147 - 0.483
TA-DistMult [13] 0.230 0.130 - 0.461
TeRo [39] 0.299 0.198 0.329 0.507
ATiSE [38] 0.252 0.148 0.288 0.462
TGeomE+ [40] 0.333 0.232 0.361 0.546
HGE [26] 0.290 0.176 0.323 0.514
ECEformer [11] 0.262 0.159 0.255 0.462
UniHR 0.333 0.240 0.367 0.527
Link Prediction on TKG.
As shown in Table 4, we achieve competitive results on the wikidata12k, even surpassing TGeomE+ by 3.4% on Hits@1 and 1.7% on Hits@3. However, existing temporal knowledge graph embedding (KGE) methods, such as TGeomE+, often employ time-aware decoders, which are challenging to generalize to other types of KGs. In contrast, our approach efficiently encodes timestamps as atomic nodes only during initialization and learns temporal information through message passing on graph structure, demonstrating that graph structure information is also beneficial for temporal knowledge graphs, highlighting the effectiveness of our UniHR.
Table 5: Ablation studies on triple prediction task.
Variant FBH FBHE DBHE
MR MRR H@10 MR MRR H@10 MR MRR H@10
w/o initial $\textbf{h}_{f}$ 5.22 0.909 0.980 5.86 0.767 0.885 2.56 0.794 0.978
w/o $\textbf{W}_{r}$ 3.06 0.944 0.992 5.98 0.792 0.885 2.30 0.850 0.976
w/o intra-fact MP 3.97 0.897 0.972 5.26 0.754 0.883 2.02 0.842 0.983
w/o $\omega_{\tau(r)}$ 2.70 0.934 0.992 5.46 0.782 0.888 2.69 0.810 0.973
w/o $\textbf{W}_{\lambda(r)}$ 2.50 0.941 0.992 5.69 0.778 0.889 2.37 0.810 0.978
w/o inter-fact MP 2.61 0.913 0.991 4.56 0.776 0.887 2.11 0.827 0.986
UniHR 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
5.3 Ablation Study on HiSL
We conduct ablation experiments on triple prediction, the most relevant task to fact nodes. As shown in Table 5, all variants with certain modules or parameters removed exhibit a decrease in performance. We can conclude that intra-fact and inter-fact message passing modules both play crucial roles, allowing UniHR to more fully represent the current fact node with enhanced structural information. We also change the initialization of fact node embeddings to learnable embeddings (w/o initial $\textbf{h}_{f}$), and results indicate that initializing fact node representations is essential. It highlights key information in the facts, mitigating the noise introduced by excessive neighbors.
5.4 Potential of Unified Representation
Table 6: Results on hyper-relational TKG datasets.
Model WiKi-hy YAGO-hy
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
TeRo [39] 0.572 0.473 0.640 0.727 0.760 0.720 0.782 0.822
HGE [26] 0.602 0.507 0.666 0.765 0.790 0.760 0.814 0.837
StarE [12] 0.565 0.491 0.599 0.703 0.765 0.737 0.776 0.820
GRAN [34] 0.661 0.610 0.679 0.750 0.808 0.789 0.817 0.842
HyNT [6] 0.537 0.444 0.587 0.723 0.763 0.724 0.787 0.836
HypeTKG [10] 0.687 0.633 0.710 0.789 0.832 0.817 0.838 0.857
UniHR 0.681 0.618 0.717 0.774 0.823 0.801 0.839 0.859
Generalize to Hyper-relational TKGs.
Due to the unity of representation, UniHR can easily generalize to compositional types of KGs, like hyper-relational TKGs (HTKGs) [10] that integrate HKGs and TKGs. Every H-fact in the HTKG is coupled with a timestamp explicitly indicating its time validity. From results in Table 6, UniHR offers a significant performance improvement in link prediction tasks on hyper-relational TKGs compared to TKG-specific and HKG-specific models, demonstrating a good collaborative modeling ability for auxiliary key-value pairs and timestamps. Our UniHR even achieves competitive performance with the specialized hyper-relational TKG model HypeTKG without complex module stacking.
Table 7: Results of separate and joint training on NKG.
FBH DBHE FBH DBHE
Model MRR H@10 MRR H@10 MRR H@10 MRR H@10
Base link prediction Triple prediction
UniHR 0.401 0.619 0.296 0.448 0.946 0.993 0.862 0.987
UniHRJoint 0.402 0.632 0.298 0.452 0.948 0.993 0.857 0.989
Joint Learning on Different Tasks of KGs.
For link prediction on NKGs, the two subtasks, namely base link prediction and triple prediction, share the same graph during the message-passing phase under our unified representation form. Therefore, we attempt joint training on the two tasks using the NKG dataset, as shown in Table 7. We observe that the results of joint training are generally superior to those of separate training, indicating that nested facts and atomic facts can mutually enhance and supplement the semantics.
Table 8: Results of separate training and joint training on the HKG and TKG dataset, where identical entities and relations share the same embeddings. WikiPeople- and wikidata12k- represent the filtered test sets.
WikiPeople- wikidata12k-
Model subject/object all entities subject/object
MR MRR H@1 H@3 H@10 MR MRR H@1 H@3 H@10 MR MRR H@1 H@3 H@10
UniHR 835.8 0.486 0.412 0.528 0.617 829.0 0.488 0.414 0.531 0.620 818.7 0.314 0.220 0.345 0.509
UniHRJoint 692.7 0.488 0.409 0.533 0.629 686.5 0.490 0.414 0.536 0.632 489.5 0.315 0.222 0.346 0.498
Joint Learning on Different Types of KGs.
We suppose that unified representation makes it possible to develop pre-trained models that integrate multiple types of KGs. To explore its potential, we conduct joint learning on different types of KGs. Therefore, we construct a hybrid dataset called wikimix which includes two subsets of Wikidata [33], namely HKG dataset WikiPeople and TKG dataset wikidata12k, which encompass 3547 identical entities and 18 identical relations. Due to the different types of facts, there are no identical facts in these two subsets. To further prevent data leakage, we filter out 537 entries from the wikidata12k test set whose main triples appear in the H-Facts of WikiPeople train set, and 384 entries from the WikiPeople test set whose main triples appear in wikidata12k train set. Statics of wikimix are given in Table 1.
From the results in Table 8, it is evident that joint learning outperforms separate learning across most metrics. Notably, there are improvements of 17.1% and 40.2% in MR metric on wikipeople- and wikidata12k- datasets, respectively. This indicates that more complex structural interactions and diverse types of training data are beneficial. Moreover, our UniHR demonstrates good scalability and effectiveness in integrating multiple types of KGs.
Table 9: Number of training parameters and times.
Dataset Model Training Params Training Times
WD50K StarE [12] 10.81M 48h+
HyNT [6] 29.61M 22.3h
HAHE [23] 42.35M 29.2h
UniHR 10.55M 23.4h
wikidata12k ATiSE [38] 31.46M 5.8h
HGE [26] 3.44M -
ECEformer [11] 11.99M 5.6h
UniHR 3.68M 3.6h
YAGO-hy HypeTKG [10] 10.83M 37.5h
UniHR 4.37M 17.6h
5.5 Efficiency Analysis
Compared to existing state-of-the-art methods, UniHR does not significantly increase the overhead in terms of model training parameters and runtime as shown in Figure 9. The embeddings of relation nodes and fact nodes are computed using a simple MLP based on atomic nodes and relations, which does not significantly increase the number of training parameters. For runtime, we do not use the whole graph during all message passing phases and adopt a dropout strategy to prevent overfitting, which can effectively enhance training efficiency. Therefore, our UniHR achieves a superior trade-off between model effectiveness and efficiency.
6 Conclusion and Future Work
In this paper, we propose UniHR, a unified hierarchical knowledge graph representation learning framework consisting of a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module. The HiDR form unifies the hyper-relational facts, nested facts and temporal facts into the form of triples, overcoming the limitations of customized encoders for different forms of facts. Moreover, HiSL captures local semantic information within facts and global structural information between facts. Our UniHR achieves the best or competitive performance across five types of KGs, and highlight the strong potential of unified representations.
In the future, we believe that the unification of knowledge graph representation learning methods is a developing trend that makes it possible to develop unified pre-trained models based on multiple types of KGs. Moreover, we will investigate how to generalize our HiDR form to more complex tasks such as inductive reasoning [30] and multi-modal scenarios, etc.
Supplemental Material Statement:
The dataset, code, and hyperparameter details are available at https://github.com/Lza12a/UniHR.
Appendix
Appendix 0.A Related Works
0.A.1 Link Prediction on Triple-based KGs
Most existing techniques in KG representation learning are proposed for triple-based KGs. Among these techniques, knowledge graph embedding (KGE) models [4, 29] have received extensive attention due to their effectiveness and simplicity. The idea is to project entities and relations in the KG to low-dimensional vector spaces, utilizing KGE scoring functions to measure the plausibility of triples in the embedding space. Typical methods include TransE [4], RotatE [29], and ConvE [9].
Depending on the KGE model alone has limitation of capturing complex graph structures, whereas augmenting global structural information with a graph neural network (GNN) [31, 25, 41] proves to be an effective approach for enhancement. The paradigm of combining GNN as encoder with KGE scoring function as decoder helps to enhance the performance of KGE scoring function. These GNN methods design elaborate message passing mechanisms to capture the global structural features. Typically, CompGCN [31] aggregates the joint embedding of entities and relations in the neighborhood via a parameter-efficient way and MA-GNN [41] learns global-local structural information based on multi-attention. These methods achieve impressive results on triple-based KGs but are hard to generalize to beyond-triple KGs.
Table 10: Base link prediction on FBHE, FBH and DBHE. All baselines’ results are taken from [37]. The best results among all models are written bold, while the second are underlined. The $s_{f}$ and $s_{h}$ denote $\left(f,has\;head\;entity,h\right)\ \left(f,has\;tail\;entity,t\right)$ and $\left(h,r,t\right)$ two types of scoring method respectively. For BiVE [7] and NestE [37], we pick their variants with best performance.
Model FBHE/FBH DBHE
MRR Hits@10 MRR Hits@10
QuatE 0.354 0.581 0.264 0.440
BiQUE 0.356 0.583 0.274 0.446
Neural-LP 0.315 0.486 0.233 0.357
DRUM 0.317 0.490 0.237 0.359
AnyBURL 0.310 0.526 0.220 0.364
BiVE 0.370 0.607 0.274 0.422
NestE 0.371 0.608 0.289 0.443
UniHR + ConvE $s_{h}$ 0.397 0.622 0.289 0.443
UniHR + ConvE $s_{f}$ 0.375 0.596 0.307 0.471
UniHR + Transformer 0.401 0.619 0.296 0.448
Appendix 0.B Decoder Analysis
To explore the effectiveness of our UniHR encoding further, we pair UniHR with different decoders and evaluated them on triple prediction task. In addition to the previously mentioned unified framework UniHR + Transformer, we also experiment on UniHR + ConvE with two scoring strategies. The ConvE [9] is the decoder customized for triples and its scoring function is $vec\left(\sigma\left(\left[\mathbf{\tilde{h}}_{h};\mathbf{\tilde{e}}_{r}\right]*\psi\right)\right)$, where $\mathbf{\tilde{h}}_{h}$ and $\mathbf{\tilde{e}}_{r}$ represent reshaped 2D embeddings of head entity $h$ and relation $r$, and * is a convolution operator. The $vec\left(\cdot\right)$ and $\psi$ are denoted as the vectorization function and a set of convolution kernels.
Due to our special representation, there exists two scoring methods for atomic triples, thus we present the base link prediction results separately for each scoring method. The $s_{f}$ represents scoring triples $\left(f,has\;head\;entity,h\right)$and $\left(f,has\;tail\;entity,t\right)$, and $s_{t}$ represents scoring $\left(h,r,t\right)$. The performance of base link prediction is shown in Table 10. Notably, FBH and FBHE share identical atomic facts, resulting in the same performance. It can be observed that regardless of the scoring method employed, we both achieve competitive performance, especially with scoring $\left(h,r,t\right)$ on FBH and scoring $\left(f,has\;head\;entity,h\right)\,\left(f,has\;tail\;entity,t\right)$ on DBHE. We attribute the differences in performance under different scoring methods to dataset characteristics. The DBHE dataset is relatively smaller, and scoring method $s_{f}$ effectively alleviates overfitting problem. Conversely, for larger datasets FBH, scoring based on $\left(h,r,t\right)$ minimizes information loss.
Table 11: Triple prediction on FBHE, FBH and DBHE. All baselines’ results are taken from [37]. The best results among all models are written bold. For BiVE [7] and NestE [37], we pick their variants with best performance.
Model FBH FBHE DBHE
MR MRR Hits@10 MR MRR Hits@10 MR MRR Hits@10
QuatE [45] 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
BiQUE [16] 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
Neural-LP [42] 115016.6 0.070 0.073 90000.4 0.238 0.274 21130.5 0.170 0.209
DRUM [28] 115016.6 0.069 0.073 90000.3 0.261 0.274 21130.5 0.166 0.209
AnyBURL [24] 108079.6 0.096 0.108 83136.8 0.191 0.252 20530.8 0.177 0.214
BiVE [7] 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
NestE [37] 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
UniHR + Transformer 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
UniHR + ConvE 3.00 0.900 0.983 6.27 0.909 0.986 2.06 0.876 0.978
UniHR + ConvE w/o $\textbf{h}_{f}$ 4.39 0.887 0.979 10.10 0.865 0.970 2.76 0.798 0.961
UniHR + ConvE w/o intra-fact 6.54 0.859 0.959 18.10 0.871 0.968 5.82 0.665 0.900
UniHR + ConvE w/o inter-fact 12.56 0.864 0.961 20.56 0.864 0.966 10.75 0.764 0.951
Table 11 shows the results of triple prediction on three benchmark datasets. Among all baselines, Quate, Bique, Neural-LP, Drum, and AnyBURL struggle to model the mapping relationship between atomic facts and nested facts. Furthermore, prior works [7] do not guarantee that all atomic facts in the nested fact test set are present in the training set as entities, which shifts the problem from a transductive setting to an inductive setting, leading to significant performance gaps between these baselines. On most metrics, our method outperforms BiVE and NestE which are specifically modeled for nested facts. Notably, NestE fully preserves the semantics of atomic facts. However, on the FBHE dataset, UniHR + ConvE achieves an improvement of 0.58 (6.4%) points in MRR and 0.24 (2.4%) points in Hits@10 compared to the state-of-the-art model NestE and the second-best performance after UniHR + Transformer on the FBH and DBHE datasets, demonstrating UniHR’s powerful graph structure encoding capabilities. We also carry out ablation experiments on UniHR + ConvE as shown in Table 11. Performance declines are observed after removing any part of the HiSL module, showing the significance of HiSL for hierarchical encoding.
Table 12: The major hyperparameters of our approach for all link prediction tasks.
Hyperparameter WikiPeople WD50K wikidata12k FBHEbase FBHbase DBHEbase FBHEtriple FBHtriple DBHEtriple
batch_size 2048 2048 2048 2048 2048 2048 2048 2048 2048
embedding dim 200 200 200 200 200 200 200 200 200
hidden dim 200 200 200 200 200 200 200 200 200
GNN_layer 2 2 2 2 2 2 2 2 2
GNN_intra-fact heads 4 4 4 4 4 4 4 4 4
GNN_intra-fact dropout 0.1 0.1 0.1 0.3 0.1 0.3 0.2 0.1 0.1
GNN_inter-fact activation tanh tanh tanh tanh tanh tanh tanh tanh tanh
GNN_dropout 0.1 0.1 0.1 0.3 0.1 0.3 0.2 0.1 0.1
transformer layers 2 2 2 2 2 2 2 2 2
transformer heads 4 4 4 4 4 4 4 4 4
transfomer activation gelu gelu gelu gelu gelu gelu gelu gelu gelu
decoder dropout 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1
soft label for entity 0.2 0.2 0.4 0.2 0.2 0.3 0.2 0.2 0.2
soft label for relation 0.1 0.1 0.3 0.2 0.2 0.3 0.2 0.2 0.2
weight_decay 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.01
learning rate 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4
Appendix 0.C Hyperparameter Settings
Here, we show the hyperparameter details for each link prediction task. To be specific, we tune the learning rate using the range $\left\{0.0001,0.0005,0.001\right\}$, the embedding dim using the range $\left\{50,100,200,400\right\}$, the GNN layer using the range $\left\{1,2,3\right\}$ and dropout using the range $\left\{0.1,0.2,0.3,0.4\right\}$. Additionally, we use smoothing label in the training phase from range $\left\{0.1,0.2,0.3\right\}$. The best hyperparameters obtained from the experiments are presented in Table 12.